OpenStudy (anonymous):

x/6 - 5 = 13 x= @sammixboo

sammixboo (sammixboo):

OKie dokie! Ofc x = me! Always! Just kidding. Onto the real problem First you must add 5 to both sides of the equation, sicne you are subtracting 5 in the original problem. x/6 - 5 + 5 = 13 + 5 x/6 = 13 + 5 So what is 13 + 5

OpenStudy (shadowlegendx):

\[\frac{ x }{ 6 } - 5 = 13\]

OpenStudy (anonymous):

haha and 18

sammixboo (sammixboo):

Okay so now plug is 18 where 13 + 5 was. x/6 = 18 Now we must make x on its own! To do this you multiply 6 on both sides of the equation. x/6 * 6 = 18 * 6 x = 18 * 6 So what is 18 * 6

OpenStudy (anonymous):

108

sammixboo (sammixboo):

Good job so now we know x = 108

sammixboo (sammixboo):

108/6 - 5 = 13 18 - 5 = 13 13 = 13

sammixboo (sammixboo):

Shad you stalker =) ;)

OpenStudy (shadowlegendx):

hehe c;